OpenStudy (anonymous):

The function f is continuous on the interval [2, 10] with some of its values given in the table below. Use a trapezoidal approximation with 4 trapezoids to approximate the integral from 2 to 10 of f of x dx.

OpenStudy (anonymous):

https://gyazo.com/beafcad65c1c6321f2fc578f0f5b7583

OpenStudy (anonymous):

@jim_thompson5910 Only person who seems to know calc

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

Just started the problem.

OpenStudy (anonymous):

I'm currently graphing online

OpenStudy (anonymous):

Just decided that my graphing calculator would be a more efficient solution.

jimthompson5910 (jim_thompson5910):

I agree. So you are able to plot the points and calculate the area of each trapezoid?

OpenStudy (anonymous):

Plotted the points, kind of stuck on the trapezoid part.

OpenStudy (anonymous):

What trapezoids are they referring to?

jimthompson5910 (jim_thompson5910):

does this graph help at all (see attached) ?

jimthompson5910 (jim_thompson5910):

here's a bigger version

OpenStudy (anonymous):

Yes, the first one seems to be a triangle though

jimthompson5910 (jim_thompson5910):

well if one parallel side of a trapezoid is 0, then we have a triangle

jimthompson5910 (jim_thompson5910):

so you can either use the triangle area formula or the trapezoid area formula (with one base being 0)

OpenStudy (anonymous):

Got 2.4 for the first one.

OpenStudy (anonymous):

4.7 for the second one.

jimthompson5910 (jim_thompson5910):

incorrect

OpenStudy (anonymous):

Seriously?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Damn calculator failed me, now I have to manually solve

jimthompson5910 (jim_thompson5910):

maybe you typed in the wrong thing? post what you typed in so I can see

OpenStudy (anonymous):

I went to STAT and put in the table above X under L1 and Y under L2

OpenStudy (anonymous):

I then went to STAT CALC and calculated a quadratic formula almost equal to the points in stat.

jimthompson5910 (jim_thompson5910):

why not just compute the area of each trapezoid using A = h*(b1+b2)/2

jimthompson5910 (jim_thompson5910):

refer to my image attachment http://assets.openstudy.com/updates/attachments/563abfcfe4b031eb5b29ec94-jim_thompson5910-1446692170541-graph3.png

OpenStudy (anonymous):

Thought I could use a shortcut.

OpenStudy (anonymous):

Which failed. So I'll do that now.

OpenStudy (anonymous):

First 2*4*1/2 = 4 Second 3*(3+8)/2 = 33/2 Third 2*(8+15)/2 = 23 Fourth (15+17)/2 = 16

OpenStudy (anonymous):

Total 59.5

OpenStudy (anonymous):

Right?

jimthompson5910 (jim_thompson5910):

`First 2*4*1/2 = 4` .... disagree; you have the wrong numbers `Second 3*(3+8)/2 = 33/2` agree `Third 2*(8+15)/2 = 23` agree `Fourth (15+17)/2 = 16` disagree; should be 15 and 18 (not 15 and 17)

OpenStudy (anonymous):

First 2*3*1/3= 3 Fourth (15+18)/2= 33/2

OpenStudy (anonymous):

Final is 59.

jimthompson5910 (jim_thompson5910):

`First 2*3*1/3= 3` agree `Fourth (15+18)/2= 33/2` agree `Final is 59.` agree

OpenStudy (anonymous):

I have a similar one but it asks me to find f'(9) should I make a new post?

OpenStudy (anonymous):

And as always, thanks.

jimthompson5910 (jim_thompson5910):

`should I make a new post?` yes please